What is the resonance frequency? 35 Hz 30 Hz, 26
Answer #1

32 . # 31 is missing. so cannot calculate resonant frequency

33.f = 10 cm
Do = 20 cm

1/f = 1/Do + 1/Di

1/Di = 1/f - 1/Do = 1/10 cm - 1/20 cm

so Di= 20 cm

34. d is answer